OpenStudy (pokemon23):

6d-12-d=10d+53 6d-12-d=10d+53 @Mathematics

OpenStudy (saifoo.khan):

d = -13

jimthompson5910 (jim_thompson5910):

6d-12-d=10d+53 5d-12=10d+53 5d=10d+53+12 5d-10d=53+12 -5d=53+12 -5d=65 d=(65)/(-5) d=-13